OpenStudy (anonymous):

Simplify. Variables represent non-negative numbers 3radical64x^7y^10=

OpenStudy (mertsj):

\[3\sqrt{64x^7y^{10}}\] \[3\sqrt{8^2x^6xy ^{10}}=3(8)(x^3)(y^5)\sqrt{x}=24x^3y^5\sqrt{x}\]